OpenStudy (anonymous):

oh my!? please help Find the largest (most positive) value of x satisfying: 3 x2 - 16 x + 16 = 0

OpenStudy (zzr0ck3r):

hint: 3x^2 - 16x + 16 = 0 is the same thing as 3x^2 + 12x + 4x + 16 = 0 now factor out some things.

OpenStudy (anonymous):

3 and 4 ? I havent done this in five years

OpenStudy (anonymous):

i can only have one answer

OpenStudy (anonymous):

of x

OpenStudy (anonymous):

just most positive value of x

OpenStudy (chihiroasleaf):

just find the solution, then pick the bigger one

OpenStudy (chihiroasleaf):

first, factor the quadratic equation, continue from the equation that @zzr0ck3r give then, you'll have two values of x, pick the bigger.. can you factor the equation?

OpenStudy (anonymous):

i have no clue how to factor isnt there like some kind of saying that helps me

OpenStudy (anonymous):

thank you for your help

OpenStudy (anonymous):

would this be 16 then? cus (x-1)(x+16)=0